[figure: 9.7 cm]

ON THE TEMPERATURE AND CONCENTRATION DRIVEN CRITICAL ELECTRON REDISTRIBUTION IN MAGNETIC

AND SUPERCONDUCTING ALLOYS

M. Matlak^{a} and M. Pietruszka^{b}

aInstitute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland bFaculty of Biology, University of Silesia, Jagiellońska 32, 40-032 Katowice, Poland

We consider a model describing magnetic and superconductors alloy Re^{(1)}_{1−x}Re^{(2)}_{x}X. We build a phase diagram of the system displaying the dependence of the critical temperatures versus concentration x. We investigate the temperature dependence of the chemical potential and average occupation numbers for a given concentration x. We find that these quantities exhibit kinks at all critical temperatures of the system. Moreover, plotting these quantities against the concentration x for a given temperatures, we also observe kinks connected with all critical concentrations of the system. This novel effect we call temperature and concentration driven critical electron redistribution and we expect experimental confirmation of this new effect.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : M. Matlak

Address for correspondence : Institute of Physics, University of Silesia, Uniwersytecka 4, PL-40007 Katowice, Poland

Email address : mpietru@us.edu.pl